### v1

Abstract.
This paper focuses on the opinion dynamics under the influence of manipulative agents. This type of agents is characterized by the fact that their opinions follow a trajectory that does not respond to the dynamics of the model, although it does influence the rest of the normal agents. Simulation has been implemented to study how one manipulative group modifies the natural dynamics of some opinion models of bounded confidence. It is studied what strategies based on the number of manipulative agents and their common opinion trajectory can be carried out by a manipulative group to influence normal agents and attract them to their opinions. In certain weighted models, some effects are observed in which normal agents move in the opposite direction to the manipulator group. Moreover, the conditions which ensure the influence of a manipulative group on a group of normal agents over time are also established for the Hegselmann-Krause model.
Key words and phrases:
Opinion dynamics, bounded confidence, manipulative groups, Hegselmann-Krause, Deffuant-Weisbuch, Weighted models
1. Introduction
The opinion dynamics is a research topic that has been quite active for several decades from various points of view: economical [6] [16], political [1] [13], mathematical [10] [2], sociological [17] [11],… Their study attempt to explain complex human behaviours when it comes to shaping individual and/or group opinions. To this end, different types of mathematical models [3] have been developed which, by their nature, can only represent very specific aspects of the mechanisms involved in the formation of opinion. Such models usually implement rules to describe the interactions between the agents involved, which may correspond to individuals, advertising messages, bots on social networks, or groups of any of these. In addition, the opinions of the agents can obey the general rules of each model or particular rules that only affect them, even if the rest of the agents interact with them following the general rules.
This can be seen most clearly when you consider that such special agents are charismatic leaders, advertising agents, or even fake news on social media who try to convince normal agents with their opinions or messages that are usually constant over time. There may also be situations where an agent wants to change their mind drastically without losing influence over normal agents who already rely on it. To do this, it will gradually modify its opinion until it arrives at the new and intended one.
For example, politicians who want to carry out a totally different policy than the one they have pursued so far, risk losing their voters. To avoid this, they can change their message over time, convincing their voters of the need for small changes, without ever discovering their ultimate intentions. After that period of time, the sum of those small changes is the big change of opinion desired to be able to propose their new policy. This example also illustrates a way of making policies acceptable that were once radical or unacceptable, thus recalling the Overton Window theory. This theory posits that policies acceptable (debatable) by society are within a window (range) and in order for society to accept other radical policies it is first necessary to widen or displace this window until society deems them acceptable. For all these reasons, it is interesting to study how the opinion dynamics evolve when in a society there are this type of special agents, which we will call manipulators, as well as strategies that they can carry out to achieve their objectives.
In this paper we will study the influence of groups of manipulative agents on discrete-time bounded confidence models such as the Deffuant–Weisbuch [4] and Hegselmann–Krause [7] models as well as the Weighted Bounded Confidence models developed in [15], [12]. We will consider the number of manipulators and the rate of change of their opinions as the fundamental variables on which our study will be based. The models studied also have dynamics that lead to different types of final distributions: a single central cluster (consensus), multiple clusters distributed throughout the opinion space, two clusters at the extremes (polarization) or even oscillating opinions in a certain range of the opinion space. For this reason, there is no general definition of the Overton Window that can be suitable for all types of final distribution. Thus, we have chosen to measure only one central point (the mean) and dispersion (the standard deviation), although the latter value does not provide much information in the polarized distributions.
The article is organized as follows. Section 2 is devoted to introduce the definitions of the different types of agents as well as the models and their dynamics. Some simulations under different parameters are presented, which serve as examples of the already known dynamics under the influence of only normal agents or a group of stubborn agents. In section 3, the dynamics of the different models under the influence of manipulators are studied. Examples are provided to illustrate the different behaviours of opinions. These examples are used to interpret the variation of the statistical values of the final distribution of opinions obtained under the action of the manipulative group. In addition, for the Hegselmann–Krause model, the conditions that ensure the interaction between one normal group and another manipulative are established based on their number of agents and the speed at which the manipulators change their minds. Finally, in section 4, the results obtained and the work that can be carried out in future articles are discussed.
2. Models of opinion dynamics under bounded confidence
In this section we will introduce the agent–based mathematical models of bounded confidence and the notation that we will use in the study of the opinion dynamics of the subsequent sections.
Let us consider a discrete agent space $P=\{p_{i}\}_{i=1}^{N}$. Each agent $p_{i}\in P$ is a subject having an opinion about some issue. The opinions can be parametrized by some real number $x\in\mathbb{R}$ (that can be thought as the amount of resources, capital…, that the agent is willing to invest or increase from a budget), in such a way that $x(p_{i})=x_{i}$. We can consider that these values are all contained in an interval $[a,b]\subset\mathbb{R}$ called the opinion space, which can be interpreted as the range of all possible opinions. The opinion of each agent can vary along the discrete time $t=0,1,2,\ldots$, so we can denote by $x_{i}(t)$ the opinion of the agent $p_{i}$ at time $t$. The main feature of bounded confidence models, or BC–models for short (see [7]), is that each agent can only be influenced by agents with similar opinion. Indeed, for the agent $p_{i}\in P$ and for some value $\varepsilon_{i}>0$, called confidence threshold, we can denote the set of confidence of $p_{i}$ by $J_{i}(t)\in\mathbb{R}$ which can be defined by
(2.1) $J_{i}(t)=\left\{k\in\{1,\ldots,N\}:|x_{i}(t)-x_{k}(t)|\leq\varepsilon_{i}\right\}$
and it is made up of the indices of agents who have their opinion at a distance no higher than the confidence threshold $\varepsilon_{i}$ of the agent $p_{i}$. The number of agents in $J_{i}(t)$ will be denoted by #$\#\left(J_{i}(t)\right)$. Observe that, #$\#\left(J_{i}(t)\right)>0$ because $i\in J_{i}(t)$ for all $i=1,\ldots,N$. The dynamics of the opinion distribution $\{x_{i}(t)\}_{i=1}^{N}$ is described, for each $x_{i}$, by an iteration such that
(2.2) $x_{i}(t+1)=x_{i}(t)+F_{i}\left(\mathbf{x}(t),t\right)$
where the functions $F_{i}$ characterize the model defining the variation of the opinion of $p_{i}$ at the discrete time $t$ and where we have denoted $\mathbf{x}(t)=\left(x_{1}(t),\ldots,x_{N}(t)\right)\in\mathbb{R}^{N}$.
There are many ways to modify these models to adapt them to different cases. For example, the number of modified opinions at every iteration $t$ leads to different models, so we will say that a model is simultaneous if the iteration given in (2.2) is applied to all agents in order to modify their opinions for each $t$. On the other hand, we will call a model alternating if, for each $t$, iteration (2.2) is applied to a single agent. It is also possible to consider the existence of repulsive opinions, that is, an opinion $x$ is repulsive for the agent $p_{i}$ if $p_{i}$ modifies his opinion $x_{i}$ by moving away from $x$.
Another important modification in the models is setting up some types of agents which do not follow the dynamics of the rest of agents. In [8], [12], several types of agents are defined for the BC–model depending on where their opinions are situated in the opinion space, whether or not their opinions evolve under the dynamics equation (2.3) and their influence over the rest of agents. So, we can define the following types of agents:
  • A normal agent is that whose opinion evolves as $x_{i}$ under the dynamics equation (2.2). They set the general type of agents and influence each other in a homogeneous way.
  • A stubborn agent is that whose opinion do not change over time. The opinions of these agents influence others but are not influenced by the rest. So, if $p_{j}$ is a stubborn agent, then $x_{j}(t+1)=x_{j}(t)$ for all $t=0,1,2,\ldots$
  • A radical agent is a stubborn agent whose opinion is close to the boundary of the opinion space $[a,b]\subset\mathbb{R}$, that is, if $p_{j}$ is a radical agent, then $x_{j}(t)=c_{j}$ constant such that $c_{j}\approx a$ or $c_{j}\approx b$ for all $t\geq 0$.
  • A charismatic agent is a stubborn agent who is not radical.
If there are $1<K\in\mathbb{N}$ stubborn agents with the same opinion in the BC–model, they can be considered as a group of agents whose influence on the neighbouring agents is equivalent to the one of an only agent with the corresponding weighted opinion. So, radical and charismatic groups are treated in [8]. Such groups can be seen as an information campaign rather than a group of actors, so they influence society but are not counted as part of it. Notice that the distinction between radical and charismatics agents is not clearly objective because we have not quantified how close to the boundary of the opinion space $[a,b]$ must be an agent to be radical.
Let us introduce a new type of agents that will not follow equation (2.2).
  • A manipulative agent or manipulator is an agent $p_{i}$ whose opinion is not influenced by the rest of agents but follows a non–constant path $x_{i}(t)=f(t)\in[a,b]\in\mathbb{R}$, where $f$ is a function with values in the opinion space $[a,b]\in\mathbb{R}$.
Manipulative agents can be thought as a generalization of stubborn agents allowing them to change their opinions without being influenced by the rest of society. This change of opinion may aim to drag the opinion of normal agents close to a limit cluster that, a priori, was not expected to reach.
In this article, we will study some variants of the classic models of Deffuant–Weisbuch [4] and Hegselmann–Krause [7] as well as the models introduced by Toccaceli et al. [15]. The first two are simple and typical models based on an iteration such that
(2.3) $x_{i}(t+1)=\sum_{k\in\Omega_{i}(t)}w_{ik}x_{k}(t)$
corresponding to the weighted arithmetical mean value of the opinions of the agents in a subset $\Omega_{i}(t)\subset J_{i}(t)$ of the set of confidence of $x_{i}(t)$ where $w=(w_{ij})$ is a matrix such that $w_{ij}\in[0,1]$ and $\sum_{k}w_{ik}=1$. In this case, the choosing of the set $\Omega_{i}$ and the weights $w_{ij}$ characterize the model of opinion.
Throughout this article, we will consider the confidence thresholds of all agents to be equal, that is $\varepsilon_{i}=\varepsilon$ for all $i=1,\ldots,N$.
2.1. The Deffuant–Weisbuch model
According to equation (2.3), when for the agent $p_{i}$ we choose the set $\Omega_{i}=\{i,j\}$ where $j\in\{1,\ldots,n\}$ is a random index such that $j\neq i$ and the weight are $w_{ij}=1/2$, then we obtain the Deffuant–Weisbuch model [4], or DW–model for short, which explicitly is
(2.4) $x_{i}(t+1)=\left\{\begin{array}[]{lcl}\displaystyle{\frac{x_{i}(t)+x_{j}(t)}{2}}&&\text{ if }\left|x_{i}(t)-x_{j}(t)\right|\leq\varepsilon\\ x_{i}(t)&&\text{ otherwise }\end{array}\right.$
This model can be implemented both in simultaneous and in alternating ways. Thus, at each iteration $t$ of the simultaneous model, a random agent $p_{j}$ is chosen for every agent $p_{i}$ with $i=1,\ldots,N$ and $i\neq j$, therefore every agent can modify its opinion simultaneously at every iteration. In the alternating model, at each iteration $t$, for a random agent $p_{i}$, another random agent $p_{j}$ is chosen to carry out the iteration (2.4). So, for each $t$, only one agent can modify its opinion. Obviously, these two ways of implementation are not equivalents although the opinion dynamics are similar. The main feature in which both forms of deployment differ is the speed of convergence. Since, in the second method, each agent intervenes in the iteration, then its convergence speed is much higher.
In figure 1, the opinion dynamics for a simultaneous Deffuant–Weisbuch model is shown with an equispaced initial distribution given by $x_{i}(0)=-1+\frac{2i}{101}$ for $i=1,\ldots,100$ and three different constant confidence thresholds $\varepsilon=0.1,0.45,0.6$. The distribution of opinions is considered stable whenever the values in each cluster differ less than $10^{-3}$ units achieved in $t_{max}$ iterations. Observe that the distances among final clustered opinions are greater than $\varepsilon$.
Figure 1. Example of opinion dynamics of the simultaneous Deffuant–Weisbuch model for a initial distribution of $N=100$ equispaced points $x_{i}(0)=-1+\frac{2i}{101}\in[-1,1]$ with $\varepsilon_{i}=\varepsilon\in\{0.1,0.45,0.6\}$ constant for all $i=1,\ldots,100$.
If a distribution of opinion interacts with a group of stubborn agents with constant opinion $x(t)=c\in[a,b]$ for all $t\geq 0$, then the opinion of agents in the confidence interval of the stubborn agents converges to this group’s opinion. A simulation of this situation can be seen in figure 2 for the same initial distribution as in figure 1 and a group of 20 stubborn agents (marked by a dashed line) with constant opinion $x(t)=0.75$ for all $t\geq 0$. Observe that the stubborn group attracts the neighbouring normal agents to its opinion.
Figure 2. Example of opinion dynamics of the simultaneous Deffuant–Weisbuch model for the same parameters as in figure 1 with a group of 20 stubborn agents with opinion $x=0.75$. The opinion of the stubborn group is drawn with a dashed line.
2.2. The Hegselmann–Krause model
In the case of the Hegselmann–Krause model [7], or HK–model, for the agent $p_{i}$, we consider the set $\Omega_{i}$ to be the whole confidence set $J_{i}$ and the new opinion $x_{i}$ of the agent $p_{i}$ is just the arithmetic mean values of all agents in $J_{i}$. So, we get
(2.5) $x_{i}(t+1)=\frac{1}{\#\left(J_{i}(t)\right)}\sum_{j\in J_{i}(t)}x_{j}(t)$
regarding that $i\in J_{i}$ for all $i$ and then, if $J_{i}(t)=\{i\}$ we have that $x_{i}(t+1)=x_{i}(t)$.
Again, analogously as for DW–model, the HK–model can be implemented as simultaneous or alternating.
To illustrate the opinion dynamics in this model, in the figure 3, we consider three simulations for an equispaced initial distribution $\{x_{i}(0)\}_{i=1}^{N}\subset[-1,1]$ for $N=100$ agents, with different constant confidence thresholds $\varepsilon\in\{0.1,0.45,0.6\}$ and $T=20$ iterations. Observe that for small values of $\varepsilon$ clusters appear at the final distribution $\{x_{i}(T)\}_{i=1}^{N}$ whereas for larger values of $\varepsilon$, the number of cluster decreases to 1, that is, consensus is reached. Note also that for intermediate values of $\varepsilon$, some clusters may appear that are not stable over time and end up grouping, reaching a consensus, after a relatively large number of iterations (see case $\varepsilon=0.45$ in the centre of figure 3), see [7].
Figure 3. Opinion dynamics in the simultaneous HK–model for a initial distribution of 100 equispaced points $x_{i}(0)=-1+\frac{2i}{101}$ for $i=1,\ldots,100$. The constant values of $\varepsilon$ are specified below every figure.
Figure 4 shows the opinion dynamics for the equispaced initial distribution $\{x_{i}(0)\}_{i=1}^{N}\subset[-1,1]$ with $N=100$, plus a group of 20 additional stubborn agents with opinion $x=0.75$. It is clear that the existence of stubborn groups change the dynamical behaviour of the opinions depending, among others, on the size of the group.
Figure 4. Opinion dynamics in the simultaneous HK–model for the same initial distribution as figure 3 with one group of 20 stubborn agents each with opinion at $x=0.75$ (marked with a dashed line).
2.3. The Weighted HK–models
It is possible to modify the BC–model assuming that the opinions change according to a weight function of the neighbouring opinions [5] in the form (2.2). In order to build an agent–based model to study the influence of fake news on society, Toccaceli et al. [15], Milli [12] define different models as in (2.2) that we will present below.
The Weighted Hegselmann–Krause model for group interactions is defined, in the opinion space $[-1,1]$, by the iteration
(2.6) $x_{i}(t+1)=\left\{\begin{array}[]{lcl}\displaystyle{x_{i}(t)+\frac{(1-|x_{i}(t)|)}{\#\left(\Omega_{i}(t)\right)}\sum_{j\in\Omega_{i}(t)}w_{ij}x_{j}(t)}&&\text{ if }\Omega_{i}(t)\neq\varnothing\\ x_{i}(t)&&\text{ otherwise }\end{array}\right.$
which modifies the value $x_{i}(t)$ by a weighted mean of the neighbouring opinions in $\Omega_{i}(t)=\{j:|x_{i}(t)-x_{j}(t)|\leq\varepsilon\text{ and }j\neq i\}$. The components of the matrix of weights $w=(w_{ij})$ describe the trustworthiness the agent $p_{i}$ attributes to the $p_{j}$’s opinion for $j\in\Omega_{i}(t)$ and their values will be taken as $w_{ij}\in[0,1]$.
Note that the term $\phi_{i}(t)=1-|x_{i}(t)|$ can be thought of as the probability density function $g(x)=1-|x|$ that gives weight to the interval $[-1,1]$. This makes moderate opinions (close to $x=0$) more likely to be modified and, therefore, after stabilization of opinion dynamics, most opinions are found in clusters $x=\pm 1$. Some opinions may fall outside these radical groups depending on the values of the confidence threshold $\varepsilon$ If the value of $\varepsilon$ is large enough, the final distribution is bimodal, polarized at $x=\pm 1$.
In figure 5, examples of opinion dynamics are shown for Weighted Hegselmann–Krause model for group interactions. The upper row shows a simulation for $N=100$ equispaced opinions of normal agents, confidence thresholds $\varepsilon=0.1,0.45,0.6$ and a random weight matrix $w=(w_{ij})$ but implemented in the alternating way (the opinion of a unique agent could be modified in each iteration). The lower row shows the simulation with the same parameters and weight matrix $w$ but implemented in a simultaneous way (the opinions of all agents could be modified in each iteration).
Figure 5. Opinion dynamics under the Weighted HK-model for group interactions for an equispaced initial distribution of $N=100$ agents for three values of $\varepsilon$ and a random matrix $w=(w_{ij})$. The graphs in the upper row have been implemented in a alternating way and the ones in the lower in a simultaneous way.
Note that the convergence rate for the simultaneous model is higher than for the alternating one.
Another model is the Weighted Hegselmann–Krause model for pairwise interactions. For the agent $p_{i}$, a random opinion $x_{j}(t)$ in the confidence interval is selected and the opinion of $p_{i}$ is modified by
(2.7) $x_{i}(t+1)=x_{i}(t)+\frac{x_{i}(t)+w_{ij}x_{j}(t)}{2}(1-|x_{i}(t)|)$
As a example of implementation of this model, in figure 6, simulations in the alternating (upper row) and simultaneous ways (lower row) for $N=100$ equispaced opinions of normal agents with three different confidence thresholds $\varepsilon=0.1,0.45,0.6$ and the same random weight matrix $w=(w_{ij})$ are shown.
Figure 6. Opinion dynamics under the Weighted HK model for pairwise interactions for an equispaced initial opinion distribution of $N=100$ normal agents for three values of $\varepsilon$ and a random matrix $w=(w_{ij})$. The upper row has been implemented in a alternating way and the lower one in a simultaneous way.
Also in [15], [12] there are some other models comprising mechanisms of attraction and/or repulsion. Let us consider an agent $p_{i}$ and a random opinion $x_{j}(t)$ in the distribution chosen such that $j\neq i$. The attractive mechanism modifies the opinion $x_{i}(t)$ by
(2.8) $x_{i}(t+1)=\left\{\begin{array}[]{lcl}\displaystyle{x_{i}(t)-s_{ij}(t)\frac{|x_{i}(t)+w_{ij}x_{j}(t)|}{2}(1-|x_{i}(t)|)}&&\text{ if }\left|x_{i}(t)-x_{j}(t)\right|\leq\varepsilon\\ \displaystyle{x_{i}(t)}&&\text{ if }\left|x_{i}(t)-x_{j}(t)\right|>\varepsilon\end{array}\right.$
where $s_{ij}(t)=sign(x_{i}(t)-x_{j}(t))$. In this case, if the opinion $x_{j}(t)$ is contained in the confidence interval of $x_{i}(t)$, then there is an attractive influence of $x_{j}(t)$ on the opinion $x_{i}(t)$, that is, the new opinion $x_{i}(t+1)$ moves closer to $x_{j}(t)$. We will call Attractive Weighted HK model to the iteration (2.8) corresponding to the attractive mechanism.
The repulsive mechanism, defining the Repulsive Weighted HK model, is given by
(2.9) $x_{i}(t+1)=\left\{\begin{array}[]{lcl}\displaystyle{x_{i}(t)}&&\text{ if }\left|x_{i}(t)-x_{j}(t)\right|\leq\varepsilon\\ \displaystyle{x_{i}(t)+s_{ij}(t)\frac{|x_{i}(t)+w_{ij}x_{j}(t)|}{2}(1-|x_{i}(t)|)}&&\text{ if }\left|x_{i}(t)-x_{j}(t)\right|>\varepsilon\end{array}\right.$
where, again $s_{ij}(t)=sign(x_{i}(t)-x_{j}(t))$. Now, if $x_{j}(t)$ is outside the confidence interval of $x_{i}(t)$ then $x_{i}(t+1)$ moves away from $x_{j}(t)$, so we consider that the influence of $x_{j}(t)$ on the opinion $x_{i}(t)$ is repulsive.
We can combine both mechanisms in one model called Attractive–Repulsive Weighted HK model. It is defined by
(2.10) $x_{i}(t+1)=\left\{\begin{array}[]{lcl}\displaystyle{x_{i}(t)-s_{ij}(t)\frac{|x_{i}(t)+w_{ij}x_{j}(t)|}{2}(1-|x_{i}(t)|)}&&\text{ if }\left|x_{i}(t)-x_{j}(t)\right|\leq\varepsilon\\ \displaystyle{x_{i}(t)+s_{ij}(t)\frac{|x_{i}(t)+w_{ij}x_{j}(t)|}{2}(1-|x_{i}(t)|)}&&\text{ if }\left|x_{i}(t)-x_{j}(t)\right|>\varepsilon\end{array}\right.$
Figure 7. Opinion dynamics under the Attractive, Repulsive and Combined Weighted HK model for an equispaced initial opinion distribution of $N=100$ normal agents, a threshold of confidence of $\varepsilon=0.1$ and the same random matrix $w=(w_{ij})$.
In the figure 7 and in order to compare the different dynamics, simulations of 100 iterations for the Attractive, Repulsive and combined (Attractive-Repulsive) models are shown for the same equispaced initial opinion distribution of $N=100$ normal agents with a confidence threshold of $\varepsilon=0.1$ and a random weight matrix $w=(w_{ij})$. When the value of $\varepsilon$ is small, the final state for the Attractive model consists of three big clusters and some individual opinions in between. The largest cluster has the opinion of about 45 normal agents and it is located at $x=0$. There are two smaller clusters located at $x=\pm 1$ with about 24 normal agents each one. The opinions of the rest of normal agents are isolated in between those three clusters. For Repulsive and Attractive–Repulsive models, the central opinions oscillate with no convergence.
As the value of $\varepsilon$ grows in the Attractive model, the distance between contiguous final opinions becomes larger until isolated opinions converge to one of the major clusters. In addition, for large values of $\varepsilon$, the final state can consist of a single cluster at $x=0$ or the two extreme clusters at $x=\pm 1$.
In the cases of Repulsive and the combined Attractive–Repulsive models, the growth of $\varepsilon$ shows dynamics with a temporal oscillation of some moderate opinions that converge, finally, to $x=\pm 1$ for larger $t$. For values $\varepsilon=0.6$ and larger, these models converge to $x=\pm 1$ without temporal oscillations.
Figure 8. Opinion dynamics under the Attractive–Repulsive Weighted HK model for an equispaced initial opinion distribution of $N=100$ normal agents for three values of $\varepsilon$ and a random matrix $w=(w_{ij})$. The upper row has been implemented in a alternating way and the lower one in a simultaneous way.
3. Opinion dynamics with manipulative agents
3.1. The Overton window
The dynamics of the model will change the initial opinion distribution and thus also the measurements that characterize its properties: mean, median,… and their associated deviations. These characteristics can be shifted in the opinion space, just as dispersion can increase or decrease, depending on the behaviour of the possible stubborn or manipulative agents. So, these changes in statistical values are reminiscent of the political theory of the Overton window. This theory, enunciated by J. P. Overton, can be applied to some political subjects in which opinions can be ordered in a opinion space in some interval $[a,b]\in\mathbb{R}$, as we do in this article. It establishes a spectrum of classification of opinions in which only a small range is considered acceptable. Outside that acceptable range, opinions can be classified as radical and unthinkable. Government policies will always be in the range of acceptability, otherwise the government will lose voters or popularity. In this theory, it is the opinions within the Overton window that enable political action. Therefore, the shift and expansion or contraction of the Overton window (of acceptable opinions) can cause previously accepted opinions to become radical or unthinkable and vice versa, affecting the policies that the government will be able to carry out.
Now, some questions arise. We may wonder what strategy a group of manipulative agents can carry out to move the Overton window in the direction that is favourable to their interests. Is it possible to quantify the size of the group of manipulators so that they can have enough influence to drag the opinion of normal agents to their target?
In order to describe the Overton window we can choose two characteristics: a center point (mean or median) and a range or a dispersion value, but there is no canonical way to define the limits of acceptable opinions. In this article, we can use various statistical values that may be appropriate to define them and study their dynamics for different models with manipulative agents, although we will focus mainly on the mean and standard deviation.
3.2. Including manipulative agents
Consider an agent space $P=\{p_{i}\}_{i=1}^{N}$ with distribution of opinion at time $t$ given by $\{x_{i}(t)\}_{i=1}^{N}$ and $K\in\mathbb{N}$ manipulative agents with opinions $\{f_{j}(t)\}_{j=1}^{K}$ at time $t$. Now, the extended set of confidence $\overline{J}_{i}(t)$ of the normal agent $p_{i}$ is defined by
(3.1) $\overline{J}_{i}(t)=\left\{k\in\{1,\ldots,N+K\}:|x_{i}(t)-z_{k}(t)|\leq\varepsilon_{i}\right\}$
where $z_{k}(t)\in\{x_{i}(t)\}_{i=1}^{N}\cup\{f_{j}(t)\}_{j=1}^{K}$ can be normal or manipulator and defined by
(3.2) $z_{k}(t)=\left\{\begin{array}[]{lcl}x_{k}(t)&&\text{ if }k\leq N\\ f_{k-N}(t)&&\text{ if }k>N\end{array}\right..$
Thus, the $Q_{i}(t)$ subset of neighbouring agents influencing $i$–th normal agent, such that $Q_{i}(t)\subset\overline{J}_{i}(t)$, can contain both manipulative and normal agents.
We are interested in the study of the opinion dynamics of normal agents $\{x_{i}(t)\}_{i=1}^{N}$ at time $t+1$, which will depend on its own opinion $x_{i}(t)$ at time $t$ and the opinions $z_{j}(t)\in Q_{i}(t)$ of the neighbouring agents, whether normal or manipulative. So, the iterations defined in the previous sections are still valid for calculating the opinion dynamics of the models, taking into account that the set of confidence needs to be extended with all manipulative agents.
We will assume that the manipulators are in a group organized in such a way that all their agents have the same opinion synchronized. With this, they achieve a greater influence on the normal opinions of their confidence interval because their common opinion has a weight of $K$ manipulative agents in the interval.
In addition, we will study the opinion dynamics in simultaneous models because in this way we are avoiding the random component of choosing, for each iteration $t$, a normal agent on which to apply the model. In this case, the model is applied to all normal agents in each iteration.
In simultaneous implementations, the set $Q_{i}(t)$ can be interpreted (specially in the DW–model and Weighted HK models) as the subset of agents having an effective influence on the opinion $x_{i}(t)$ in the interval of time $[t,t+1]$, so some others agents in the confidence interval can interact with the agent $p_{i}$ but they do not have real influence on it.
3.3. The HK–model with manipulators
Consider an agent space $P=\{p_{i}\}_{i=1}^{N}$ with distribution of opinion at time $t$ given by $\{x_{i}(t)\}_{i=1}^{N}$ and $m\in\mathbb{N}$ groups of stubborn or manipulative agents with opinions $\{f_{j}(t)\}_{j=1}^{m}$ at time $t$. We assume that the $j$–th group comprises of $k_{j}\in\mathbb{N}$ agents and the confidence threshold $\varepsilon>0$ is constant for all $i=1,\ldots,N$. Then, the dynamics of this model is ruled by
(3.3) $x_{i}(t+1)=\frac{1}{\#\left(\overline{J}_{i}(t)\right)}\sum_{j\in\overline{J}_{i}(t)}z_{j}(t)=\frac{1}{\#\left(J_{i}(t)\right)+\underset{j>N}{\sum_{j\in\overline{J}_{i}(t)}}k_{j}}\left[\underset{j\leq N}{\sum_{j\in\overline{J}_{i}(t)}}x_{j}(t)+\underset{j>N}{\sum_{j\in\overline{J}_{i}(t)}}k_{j}f_{j}(t)\right]$
according to (3.2), where $\overline{J}_{i}(t)$ is defined as in (3.1) and $J_{i}(t)$ as in (2.1).
In the figures 9 and 10, we illustrate the opinion dynamics of $N=100$ normal agents with an equispaced initial distribution in the opinion space $[-1,1]$ and an only group of 8 manipulators during 200 iterations with a confidence threshold $\varepsilon=0.1$. Each of the six graphics describes the effect of the group of manipulators moving theirs opinions linearly from $f(0)=-0.8$ to $f(t_{\Delta})=0.8$ for $t_{\Delta}\in\{35k\}_{k=1}^{6}$ iterations and, after, $f(t)=0.8$ for all $t>t_{\Delta}$. Observe that the value $t_{\Delta}$ corresponds to the number of iterations in which the manipulators reach their final opinion.
Figure 9. Simultaneous HK–model: Trajectories of an equispaced initial opinion distribution of 100 normal agents and a group of 8 manipulators during 200 iterations with a confidence threshold $\varepsilon=0.1$. The slopes of the increasing segment of the opinion $x=f(t)$ of manipulators correspond to $t_{\Delta}\in\{35k\}_{k=1}^{6}$ iterations to achieve the final opinion $f(t_{\Delta})=0.8$ from the initial $f(0)=-0.8$. The dashed line represents the trajectory of the opinion of the manipulative group.
Figure 9 shows that the manipulative group “pulls” the opinions of normal agents such that, for fixed confidence threshold, numbers of normal and manipulative agents, the number of clusters depends on the speed at which the opinion of manipulators changes. In this model, the group of manipulative agents is not always able to convince the normal agents in its confidence interval because the speed with which it changes its opinion is not always slow enough to attract a large number of normal agents with them, leaving the confidence interval of the group. In the figure 10, the orange bars represent graphics opinion/number of agents of the final clusters at time $T=200$. Observe that, in this example, for manipulators who change their minds at higher speeds, the most populated clusters remain at the opposite side of the final opinion of the group of manipulators and as that speed decreases there is a displacement of the clusters in the same direction as the group does. This effect will be explained in section 3.3.2.
Figure 10. Simultaneous HK–model: Clusters with their number of agents. The black arrow and dashed lines indicate the movement of the opinion of the manipulative agents corresponding to the dynamic of figure 9.
3.3.1. Dynamic with one manipulative group
It is known that, for a symmetric equispaced initial distribution $\mathbf{x}(0)=\{x_{i}(0)\}_{i=1}^{N}$, the mean and the median opinions $\overline{\mathbf{x}}$, $\mathbf{x}_{M}$ in the HK–model (2.3) without any manipulative agent in the opinion space $[-1,1]$ are $\overline{\mathbf{x}}=0$ and $\mathbf{x}_{M}=0$ (see figure 3). The deviation values of the final distribution of opinions $\mathbf{x}(T)=\{x_{i}(T)\}_{i=1}^{N}$ depends on the numbers of final clusters. To analyse the behaviour of the model with a only group of manipulators, we will set the confidence threshold to $\varepsilon=0.1$ to compare the results obtained with a final distribution with several clusters in the case of the dynamics in which only normal agents intervene.
As a first simulation, we will use the initial distribution $\mathbf{x}(0)=\{-1+\frac{2i}{101}\}_{i=1}^{100}$ of normal agents and a group of $K$ manipulators whose opinion changes linearly from $f(0)=-0.8$ to $f(t_{\Delta})=0.8$ and $f(t)=0.8$ for $t\geq t_{\Delta}$. The values of $K$ and $t_{\Delta}$ are chosen to be $K\in\{0,\ldots,30\}$ and $t_{\Delta}\in\{10+6j\}_{j=0,\ldots 30}=\{10,16,22,\ldots,190\}$. The criterion for stopping iterations is
[…]
The aim of this simulation is to observe the behaviour of the mean values of the final distribution of the normal agents as well as its dispersion measured by the standard deviation of the complete final distribution.
In the graph on the left of figure 11, we represent the mean value of the final distribution $\mathbf{x}(200)$ by a color map depending on the number $t_{\Delta}$ of iterations ($x$–axis) and the number $K$ of agents in the manipulative group ($y$–axis). We observe that, for small values of the variables, the effect of the manipulative group is not very significant (in terms of the mean value) since the mean values remain close to the mean achieved without manipulators, i.e. $\overline{\mathbf{x}}\simeq 0$, illustrated in the graph by dark blue colors.
Figure 11. Simultaneous HK–model: Mean value (left), standard deviation (centre) and number of final clusters (right) depending on the number $K$ of manipulators and the number $t_{\Delta}$ of iterations the manipulators are using to reach their final opinion $f(t_{\Delta})=0.8$ from their initial opinion $f(0)=-0.8$. The confidence threshold is fixed at $\varepsilon=0.1$ constant.
Small values of $t_{\Delta}$ imply high speeds in the change of opinion of the manipulators. When the values of $t_{\Delta}$ are increased, the opinion of this group changes more slowly and this allows the manipulators to shift the average value of the opinion of normal agents in the same direction as the opinion of the manipulative group. Bright yellow indicates mean values close to the final opinion of the manipulators.
The graph on the centre of figure 11 shows the standard deviation of the final distribution of normal agents. We can observe that, when the group of manipulators has little influence on normal agents (small values of $K$ and $t_{\Delta}$), the standard deviation of the final distribution takes values $\sigma\simeq 0.55$ slightly higher than the value without manipulators $\sigma=0.5342$ (illustrated in yellow colors). As the variables grow, it can be observed that the standard deviations decrease to $\sigma=0$ (in dark blue), indicating that consensus is reached. This consensus is achieved in the opinions close to the final opinion of the group of manipulators $f(t)=0.8$ for $t\geq t_{\Delta}$. There are no significant differences in the dispersion behaviour if we use the mean absolute deviation, so we only show the standard deviation values $\sigma$.
On the right of figure 11, we can see the number of final clusters, where dark blue represents consensus, that is, all opinions converge to only one cluster.
Figure 12. Simultaneous HK–model: Median value (left), median absolute deviation (centre) and number of iterations in which manipulators behave like stubborn agents after reaching their final opinion $f(t_{\Delta})=0.8$ until the stopping criterion is satisfied (right).
In figure 12, we can observe a similar behaviour of the median value (left) and the median absolute deviation (centre). The graph on the right shows the number of iterations required, after the group of manipulators reaches their final opinion, for the distribution to stabilize.
A complementary way to study how a manipulative group changes the dynamics of opinion is to consider the range of $\alpha$of opinions around the median value, that is, we consider the dynamics of interval $I(t)=[x_{a}(t),x_{b}(t)]$ in the opinion space such that it contains only $\alpha$of the central opinions and excludes $\frac{1}{2}(100-\alpha)$of the most extreme opinions on each side of the opinion space. In our simulation, we set $\alpha=90$% and since $N=100$ normal agents, so once the vector $\mathbf{x}(t)$ is sorted in ascending order, we get the interval $I(t)=[\mathbf{x}_{6}(t),\mathbf{x}_{95}(t)]$. In figure 13, we can observe after the last iteration $t=t_{max}$, the distances of $x_{6}$ and $x_{95}$ to the median value $\mathbf{x}_{M}$ respectively (graphs on the left and center). To study the symmetry of the interval $I$, we compute the difference $d_{-}-d_{+}$ between the distances $d_{-}=|x_{6}-\mathbf{x}_{M}|$ and $d_{+}=|x_{95}-\mathbf{x}_{M}|$. In the graph on the right, the sea green color represents the symmetry of the interval. The blue colors indicate that subinterval $[\mathbf{x}_{M},\mathbf{x}_{95}]$ is larger than the subinterval $[x_{6},\mathbf{x}_{M}]$, while the yellow colors represent the opposite case. It can be seen that, when consensus is not reached, there is a curved blue stripe in which the distance $d_{+}$ is greater than $d_{-}$, which corresponds to a final distribution of two clusters (see figure 13 right). Below that stripe, the three cases are mixed without there being clearly differentiated regions, because the median $\mathbf{x}_{M}$ jumps from one cluster to another with small variations of $t_{\Delta}$ and $K$.
Figure 13. Simultaneous HK–model: Length of the subintervals $[x_{6},\mathbf{x}_{M}]$ (left) and $[\mathbf{x}_{M,}x_{95}]$ (centre) and difference between both length (right).
Thus, it is noted that, when there is intervention of a manipulative group, the central values of the opinion (mean and median) of the normal population move in the same direction as the manipulators (from $\overline{\mathbf{x}}=0$ and $\mathbf{x}_{M}=0$ without manipulators) while the deviations decrease. This tells us that the manipulators manage to shift and narrow Overton’s window on the subject of study. The narrowing in the range of opinions is also due to the nature of the HK–model, since it tends to concentrate opinions on intermediate values.
3.3.2. Analysis of the opinion dynamics among groups of manipulators and normal groups
Notice, as shown in figures 3, 4 and 9, that clusters take shape in each confidence interval after a few iterations, so it is appropriate to analyse the interaction among groups of normal agents and groups of manipulators.
If we consider $m$ manipulative groups with opinion $\{f_{j}(t)\}$ and $K_{j}\in\mathbb{N}$ agents for $j=1,\ldots,m$ interacting with $n$ groups with $N_{s}\in\mathbb{N}$ normal agents with opinion $\{x_{i(s),s}(t)\}$ for $s=1,\ldots,n$ and $i=1,\ldots,N_{s}$ in the same confidence interval then, by iteration (3.3) and for any $\alpha=1,\ldots,n$ and $\beta=1,\ldots,N_{\alpha}$, we have
(3.4) $x_{\beta,\alpha}(t+1)=\frac{\sum_{j=1}^{m}K_{j}f_{j}(t)+\sum_{s=1}^{n}N_{s}x_{i(s),s}(t)}{\sum_{j=1}^{m}K_{j}+\sum_{s=1}^{n}N_{s}}$
and since the left hand side of the expression does not depend on $\alpha$ or $\beta$, then all normal groups reach a local consensus in the first iteration becoming an aggregated normal group of $N=\sum_{s=1}^{n}N_{s}$ agents with the same opinion. Moreover, by the linearity of the mean values, we can simplify the study of the interaction among several groups of normal agents and manipulators to the interaction of just one group of each type. The opinions of such groups will be the mean values of the opinions of all agents in each one. Note that these new groups will be “virtual” because there may not be real agents with its opinion.
So, let us consider a group of $N$ normal agents such that $x_{i}(0)=x_{0}$ for $i=1,\ldots,N$ and a group of $K$ manipulative agents such that their opinions are defined by $f(t)=f_{0}+\lambda t$ with $|f_{0}-x_{0}|<\varepsilon$ and $\lambda\in\mathbb{R}$. Since $x_{i}(0)=x_{j}(0)$ for all $i,j=1,\ldots,N$ then we can do without sub-indexes naming the opinion of normal agents by $x=x_{i}$. Then we have that opinion dynamics follows the formula
$x(t+1)=\frac{Kf(t)+Nx(t)}{K+N}$
for $t\geq 0$.
Proposition 3.1.
If $|f(t)-x(t)|\leq\varepsilon$ and $|\lambda|\leq\frac{K\varepsilon}{K+N}$ then $|f(t+1)-x(t+1)|\leq\varepsilon$.
Proof.
Assuming the hypotheses, we have
$\left|f(t+1)-x(t+1)\right|$ $=\left|\left(f(t)+\lambda\right)-\left(\frac{Kf(t)+Nx(t)}{K+N}\right)\right|=$
$=\left|\frac{(K+N)f(t)+\lambda(K+N)-Kf(t)-Nx(t)}{K+N}\right|=$
$=\left|\frac{N\left(f(t)-x(t)\right)+\lambda(K+N)}{K+N}\right|=$
$=\left|\frac{N}{K+N}\left(f(t)-x(t)\right)+\lambda\right|\leq$
$\leq\left|\frac{N}{K+N}\left(f(t)-x(t)\right)\right|+\left|\lambda\right|\leq$
$\leq\frac{N\varepsilon}{K+N}+\left|\lambda\right|\leq$
$\leq\frac{N\varepsilon}{K+N}+\frac{K\varepsilon}{K+N}=\varepsilon$
as we claimed. ∎
Proposition 3.2.
If $|f(t)-x(t)|\leq\varepsilon$ for all $t\in\{0\}\cup\mathbb{N}$, then we have
(3.5) $f(t)-x(t)=\left(\frac{N}{K+N}\right)^{t}\left(f_{0}-x_{0}\right)+\frac{\lambda(K+N)}{K}\left(1-\left(\frac{N}{K+N}\right)^{t}\right)$
for all $t\in\{0\}\cup\mathbb{N}$.
Proof.
By hypothesis, the opinion of the group of normal agents is under the influence of the group of manipulators for all $t\in\{0\}\cup\mathbb{N}$. So, we can prove the result by induction. Indeed, the formula (3.5) is trivially verified for $t=0$. Assuming equality (3.5) is true for $t$, we can write that
$f(t+1)-x(t+1)$ $=\left(f(t)+\lambda\right)-\left(\frac{Kf(t)+Nx(t)}{K+N}\right)=$
$=\frac{(K+N)f(t)+\lambda(K+N)-Kf(t)-Nx(t)}{K+N}=$
$=\frac{N\left(f(t)-x(t)\right)+\lambda(K+N)}{K+N}=$
$=\frac{N\left[\left(\frac{N}{K+N}\right)^{t}\left(f_{0}-x_{0}\right)+\frac{\lambda(K+N)}{K}\left(1-\left(\frac{N}{K+N}\right)^{t}\right)\right]+\lambda(K+N)}{K+N}=$
$=\left(\frac{N}{K+N}\right)^{t+1}\left(f_{0}-x_{0}\right)+\frac{\lambda N}{K}+\lambda-\frac{\lambda(K+N)}{K}\left(\frac{N}{K+N}\right)^{t+1}=$
$=\left(\frac{N}{K+N}\right)^{t+1}\left(f_{0}-x_{0}\right)+\frac{\lambda(K+N)}{K}\left(1-\left(\frac{N}{K+N}\right)^{t+1}\right)$
so, the formula (3.5) is also true for $t+1$ and, by the hypothesis of induction, it is satisfied for all positive integer $t\in\{0\}\cup\mathbb{N}$. ∎
A trivial consequence of the proposition 3.1 is the following corollary.
Corollary 3.3.
The following conditions
  1. (1)
    $|f_{0}-x_{0}|\leq\varepsilon$ and $|\lambda|\leq\frac{K\varepsilon}{K+N}$.
  2. (2)
    $|f(t)-x(t)|\leq\varepsilon$ for all $t\in\{0\}\cup\mathbb{N}$.
are equivalent.
Proof.
(1) $\Rightarrow$ (2). It is trivial by mathematical induction using proposition 3.1.
(2) $\Rightarrow$ (1). By equation (3.5), we have
$\left|f(t)-x(t)\right|=\left|\left(\frac{N}{K+N}\right)^{t}\left(f_{0}-x_{0}\right)+\frac{\lambda(K+N)}{K}\left(1-\left(\frac{N}{K+N}\right)^{t}\right)\right|$
and computing limits
$\lim_{t\mapsto\infty}\left|\left(\frac{N}{K+N}\right)^{t}\left(f_{0}-x_{0}\right)+\frac{\lambda(K+N)}{K}\left(1-\left(\frac{N}{K+N}\right)^{t}\right)\right|\leq\varepsilon$
we obtain
$\left|\frac{\lambda(K+N)}{K}\right|\leq\varepsilon\qquad\Longrightarrow\qquad\left|\lambda\right|\leq\frac{K\varepsilon}{K+N}$
as claimed. ∎
The condition (1) of corollary 3.3, that is,
(3.6) $\left|\lambda\right|\leq\frac{K\varepsilon}{K+N}\quad\text{ with }\left|f_{0}-x_{0}\right|\leq\varepsilon$
gives us the relationship among the quantities $K$, $N$, $\lambda$ and $\varepsilon$ to ensure that the group of manipulators influences the group of normal agents for all $t>0$. Observe that $\lambda$ describes the speed of change of the opinion of the manipulative group, so when this rate decreases, the ability of the manipulators to influence normal agents increases. Similarly, when the proportion of manipulative agents to the total number of agents (manipulators and normals) is increased, which is given by $w_{f}=\frac{K}{K+N}$, the ability of the manipulator group to influence the normal population is also increased. The same effect has the confidence threshold $\varepsilon$ so, in the HK–model, the more distrustful the population (small values of $\varepsilon$) the lower the influence of the manipulative group. If each normal agent trusts any other opinion (large values of $\varepsilon$), the manipulative group will attract the opinion of the group (or groups) of normal agents to its own, thus achieving the manipulation of the whole society.
Corollary 3.3 implies that, whenever $\left|\lambda\right|>\frac{K\varepsilon}{K+N}$ with $\left|f(t_{0})-x(t_{0})\right|\leq\varepsilon$, there will exist $T\geq 1$ such that $\left|f(t_{0}+T)-x(t_{0}+T)\right|>\varepsilon$. Then, the influence of manipulators stops at $t=t_{0}+T$ leaving the confidence interval of the normal group. This can be seen in figures 9 and 10. The manipulative group attracts normal agents into a cluster which follows the trajectory of opinion of manipulators until the number $N$ of normal agents is big enough not to verify the condition $\left|\lambda\right|\leq\frac{K\varepsilon}{K+N}$. Thus, the manipulative group leaves the confidence interval of the cluster of normal agents whose opinion is not longer modified because there is not another agent in said interval.
3.4. The DW-model with manipulative agents
Using the same notation as in previous sections, we can define the iteration for the DW–model with manipulators by
$x_{i}(t+1)=\left\{\begin{array}[]{lcl}\displaystyle{\frac{x_{i}(t)+z_{j}(t)}{2}}&&|x_{i}(t)-z_{j}(t)|\leq\varepsilon\\ x_{i}(t)&&|x_{i}(t)-z_{j}(t)|>\varepsilon\end{array}\right.$
where $z_{j}$ is the opinion of a random agent (normal or manipulator).
In figures 14 and 15, we show a simulation for the DW–model with the same parameters as in simulation of figures 9 and 10 except the number of iterations until manipulators reach their final opinion, which are chosen $t_{\Delta}\in\{120k\}_{k=1}^{6}$. Observe that we obtain a similar dynamics but with a greater fragmentation in final clusters.
Figure 14. Simultaneous DW–model: Trajectories of an equispaced initial opinion distribution of 100 normal agents and a group of 8 manipulators with a confidence threshold $\varepsilon=0.1$. The manipulative group changes its opinion lineally from $f(0)=-0.8$ to $f(t_{\Delta})=0.8$ for $t_{\Delta}\in\{120k\}_{k=1}^{6}$ iterations. The dashed line represents the trajectory of the opinion of the manipulative group.
Figure 15. Simultaneous DW–model: Clusters with their number of agents. The black arrow and dashed lines indicate the range of the opinion of the manipulative agents corresponding to the dynamics of figure 14.
Next, we will study the DW–model in a similar way as we have done in subsection 3.3.1, but in this case, due to its random component, we will perform 100 simulations for each pair of fixed values $(t_{\Delta},K)$, calculating the corresponding average of the statistical values. It should be noted that medians and median absolute deviations behave, on average, very similarly to means and standard deviations, so we will only show the latter.
The initial distribution of normal agents for all simulations is $\mathbf{x}(0)=\{-1+\frac{2i}{101}\}_{i=1}^{100}$ and, again, we consider a group of $K$ manipulators whose opinion changes linearly from $f(0)=-0.8$ to $f(t_{\Delta})=0.8$ remaining constant $f(t)=0.8$ for $t\geq t_{\Delta}$. The range of $K$ will be the same as for the HK-model: $K\in\{0,\ldots,30\}$ but the values of $t_{\Delta}$ are chosen higher since the convergence is slower. So, we will take $t_{\Delta}\in\{30+30j\}_{j=0}^{30}=\{30,60,90,\ldots,930\}$ and set the stopping criterion of iterations by making it so that, after each iteration, the values of the opinions, rounded to 3 decimal places, are clustered at a distance among clusters greater than the confidence threshold $\varepsilon=0.1$.
In figure 16, it is observed (left) that the mean values increase as the number of the iterations $t_{\Delta}$ (slow change in the opinion of manipulators) and/or the number $K$ of manipulative agents in the group increase. The brightest yellow indicates mean values close to $x=0.72$, which do not reach the value of the final opinion of the manipulative group. This is because there is a certain dispersion (center) in the final distribution as small groups appear outside the range of opinion of the manipulative group. Due to the random nature of the model, the manipulative group is not able to influence, at each iteration, all normal agents in the confidence interval. Note that the mean value of the number of clusters (right) is in the range $[2,11]$, so there is generally no consensus. In fact, in this third graph, the darkest blue represents mean values close to $c=2.5$ clusters. Notice that, in figures 14 and 15, we can observe that, for $t_{\Delta}=720$, the simulation stops with three final clusters but, due to the random character of this model, it is possible to get simulations with a lesser number of final clusters.
Figure 16. Simultaneous DW–model: Means of mean values (left), standard deviation (centre) and number of final clusters (right) after 100 simulations for every number $K$ of manipulators in the group and every number $t_{\Delta}$ of iterations, with a constant confidence threshold $\varepsilon=0.1$.
Regarding the range of $\alpha$of opinions around the median, we can see that when we take $\alpha=90$%, there is an asymmetry in the interval since the subinterval $[x_{6},\mathbf{x}_{M}]$ with values lower than the median is longer than the subinterval $[\mathbf{x}_{M},x_{95}]$ with higher values (see figure 17). But if we consider $\alpha=80$% or lesser, we do not take into account the radical agents not influenced by the manipulative group with opinions close to $x=\pm 1$, and the symmetry appears even when large number of final clusters are obtained.
Figure 17. Simultaneous DW–model: For $\alpha=90$% we represent the length of the subintervals $[x_{6},\mathbf{x}_{M}]$ (left) and $[\mathbf{x}_{M,}x_{95}]$ (centre) and difference between both length (right).
As in the case of the HK-model, in the simultaneous DW-model, the manipulative group moves the mean values of the opinion of the normal agents in the same direction as the manipulators, but whose influence is not so great as to reach consensus on the opinion final value of the manipulators. In fact, the final opinions of normal agents are grouped into a larger number of clusters.
3.5. The Weighted HK models with manipulative agents
In this section, we will consider the simultaneous Weighted HK models interacting with one group of manipulators. The iterations in these models are analogous to expressions (2.6), (2.7), (2.8), (2.9) and (2.10) except we have to substitute $x_{j}$ by $z_{j}$ where the latter will be an agent (manipulator or normal) in the confidence interval of the normal agent $x_{i}$.
3.5.1. Group and pairwise interactions models
Both the model for group and for pair interactions have a strong tendency to converge, bi-modally, to the opinions in the extremes of the range. This means that a big group of manipulators (even twice the size of the entire normal population) does not alter, at least significantly, the final distribution of opinions. In these models, convergence occurs with a small number of iterations, so if the opinion of the manipulative group opinions varies slowly, the normal agents have enough time to approach the extreme opinions so that the manipulators do not find normal agents in their confidence interval. Only when the number of manipulator is large (the same size as the normal population or larger), they are able to carry some of the few moderate opinions of normal agents to the extremal opinion nearest to the final opinion of manipulators.
3.5.2. Attractive and Repulsive Weighted HK models
In the case of the Simultaneous Attractive Weighted HK model, the opinion dynamics with a manipulative group is richer than for the previous models depending on the values of the threshold $\varepsilon$, the number of iterations $t_{\Delta}$ and the ratio between the number of manipulators in the group and the number of normal agents.
As described in section 2.3, whenever the value of $\varepsilon$ is small (see the graphic at the left of figure 7), the final distribution comprises three major clusters at $x=-1,0,+1$ with some few isolated opinions in between. So, the influence of the manipulative group provokes perturbations on the major clusters and can modify the isolated opinions even grouping them to some cluster. This behaviour is illustrated in figure 18, where a simulation is done with 100 normal opinions initially equispaced and a manipulative group of 60 agents for a confidence threshold $\varepsilon=0.1$. The number of iterations $t_{\Delta}$ to move the opinion of the manipulators from $f(0)=-0.8$ to $f(t_{\Delta})=0.8$ is chosen as $t_{\Delta}\in\{60,120,180\}$.
Figure 18. Trajectories of 100 normal opinions initially equispaced normal agents and a group of 60 manipulators according to the Simultaneous Attractive Weighted HK model. The confidence threshold is fixed to $\varepsilon=0.1$ and the iterations $t_{\Delta}\in\{60,120,180\}$. The initial opinion of manipulators is $f(0)=-0.8$ and the final opinion is $f(t_{\Delta})=0.8$. The trajectory of the opinion of manipulative group follows the dashed line.
For larger values of $\varepsilon$, the distribution converges into two clusters at $x=\pm 1$, with similar number of agents each. Adding a manipulative group, it can move the opinions of normal agents from one side of the opinion space to the other, depending on the size of the group and the trajectory of its opinion. The simulation of figure 19 is made on an equispaced initial distribution of 100 normal agents and a manipulative group of 60 agents for a confidence threshold $\varepsilon=0.6$. Observe that the opinion of manipulators, again, runs from $f(0)=-0.8$ to $f(t_{\Delta})=0.8$ in $t_{\Delta}\in\{10,20,30\}$ iterations. For the value $t_{\Delta}=10$ (left), the final distributions consist of a cluster composed of approximately half of the opinions at $x=-1$ and the rest oscillates between $x=1$ and $x=0.8$ because the manipulators become stubborn within the confidence interval of normal agents close to $x=1$. In the case of $t_{\Delta}=20$ (center), the manipulators move the opinions formerly in the cluster $x=-1$ to a new cluster at $x=0$, and the rest of opinions still oscillate between $x=1$ and $x=0.8$. Finally, for $t_{\Delta}=30$ (right), the manipulators move all the opinions to an only cluster $x=1$. In this case, the number of manipulators is lesser than the number of normal agents (the entire normal population), so the manipulative group is not strong enough to cause the opinion of normal agents to oscillate for many iterations between $x=1$ and $x=0.8$.
Figure 19. Trajectories of an equispaced initial distribution of 100 normal agents and a group of 60 manipulators according to the Simultaneous Attractive Weighted HK model. The confidence threshold is fixed to $\varepsilon=0.6$ and the iterations $t_{\Delta}\in\{10,20,30\}$. The initial opinion of manipulators is $f(0)=-0.8$ and the final opinion is $f(t_{\Delta})=0.8$. The trajectory of the opinion of manipulative group follows the dashed line.
As in section 3.3.1, we will analyse how a manipulative group with $K$ agents affects the opinion dynamics of the Simultaneous Attractive Weighted HK model. We will proceed by fixing the confidence threshold $\varepsilon$ but varying the number of iterations $t_{\Delta}\in\{5j\}_{j=1}^{30}$ and the number of manipulators in the group as $K\in\{5j\}_{j=0}^{30}$. For every pair $(t_{\Delta},K)$, we will do 100 simulations and compute the average values obtained from the mean opinions of normal agents, from the standard deviations and from the number of final opinions (maybe clusters and oscillating or isolated opinions). We have to take into account that, as seen in figures 18 and 19, we can find cases in which the opinion dynamics does not converge and there exists normal agents with oscillating opinions. So, for each simulation, we will consider the opinion distribution in two particular iterations: $t=t_{\Delta}$ where the manipulative group reaches its final opinion and $t=500$ which allows the stabilization of the distribution in case it occurs. For the values $t\in[t_{\Delta},500]$ the manipulative group becomes stubborn.
Figure 20 shows the results for fixed $\varepsilon=0.6$. The upper row corresponds with the values computed at $t=t_{\Delta}$ while the lower row shows the results at $t=500$. The comparative between these two classes of results permits to distinguish the influence of the manipulators to move the opinion of normal agents and their ability (as stubborn group) to retain the opinion of the latter. Observe that the scales for $t_{\Delta}\in[5,150]$ and $K\in[0,150]$ have been chosen larger than for the DW and HK models (cf. sections 3.3 and 3.4) due to the little influence that a small or rapidly changing manipulative group has on the normal population.
In the case of $t=t_{\Delta}$ (upper row) we can observe the following. For $K<30$ with $t_{\Delta}$ greater than $K$, the average of mean opinions is close to $x=0$ (dark blue in the left picture) with a average of standard deviations close to $\sigma=1$ (yellow in the centre) and a small number of final opinions (dark blue in the right). This implies that the opinion distribution is near a bimodal distribution $x=\pm 1$. Whenever $t_{\Delta}<20$, the average of mean opinions is also close to $x=0$ but with smaller dispersion $\sigma\simeq 0.6$ (green in the centre) and a large number of different opinions (yellow at the right picture). From these regions, through intermediate states of transition, one arrives at $t_{\Delta}\gtrsim 40$ and $K\gtrsim 40$, where the manipulators bring the opinion of the normal agents close to their own. Approximately if $K\in[50,100]$ with $K<t_{\Delta}$ (yellow area in the left), the average opinions of normal agents are close to $x=1$ since the number of manipulative agents is not large enough to retain them in their final opinion $x=0.8$. If $K>50$ and $K>t_{\Delta}$ (orange in the left picture), the manipulative group achieves that the average opinion of normal agents is around $x=0.8$, although with some dispersion $\sigma\simeq 0.3$ due to oscillating opinions.
Now, if we compare with the results after $t=500$ iterations (lower row), we can observe that the region with the average of mean opinions $x=1$ (yellow in the left picture) is clearly delimited below $K=100$. This implies that the manipulative group is able to attract normal agents to its final opinion $x=0.8$, but it cannot retain them in that same opinion. It is the fixed point $x=1$ of the model that ends up attracting them. Only when the size of the manipulative group is greater than the total population of normal agents, that is $K>100$, the opinion of the normal agents oscillates around $x=0.8$ obtaining that average value but with dispersion.
Figure 20. Simultaneous Attractive Weighted HK model. After 100 simulations for every numbers $K\in\{5j\}_{j=0}^{30}$ of manipulative agents in the group and $t_{\Delta}\in\{5j\}_{j=1}^{30}$ of iterations that the manipulators are using to vary their initial opinion $f(0)=-0.8$ to their final opinion $f(t_{\Delta})=0.8$, the pictures show the mean values of the mean opinions (left), of the standard deviation (centre) and of number of final opinions (right). The confidence threshold is fixed at $\varepsilon=0.6$ constant.
It is worth mentioning that for small confidence intervals, for example $\varepsilon=0.1$, the influence of the manipulative group at $t=t_{\Delta}$ has an opposite result to what is intended. Notice in the figure 21 that far from moving the opinions of the normal agents towards $x=0.8$, what is achieved is a slight shift in the opposite direction. The blue area in the picture on the left indicates negative values. Also observe that, for all values of $K$ and $t_{\Delta}$ in the simulation, we obtain a moderated dispersion and a moderate or high number of final opinions for $t=t_{\Delta}$.
Figure 21. Simultaneous Attractive Weighted HK model. Simulation with the same parameter as the one in figure 20 except that the confidence threshold is fixed at $\varepsilon=0.1$.
In the case of the Simultaneous Repulsive HK model, recall that it essentially has two behaviours. When the confidence threshold $\varepsilon$ is small, opinions oscillate in the central part of the opinion space so that, after some iterations, some opinions leave that central zone to converge at one of the extremes. If the confidence threshold increases, so does the number of opinions coming out the central zone, so that if $\varepsilon$ is large, the oscillations disappear and normal opinions converge to one of the extremes.
When a manipulative group acts, normal opinions tend to separate from manipulative ones when the former are outside the latter’s confidence interval. Therefore, the smaller the confidence interval, the greater the number of normal opinions that are able to drift away.
If we run a simulation with 100 normal agents, a manipulative group with $K\in\{5j\}_{j=0}^{30}$ agents who change their opinions from $x=-0.8$ to $x=0.8$ in $t_{\Delta}\in\{5i\}_{j=1}^{30}$ iterations, we obtain the graph in figures 22 and 23.
Figure 22. Simultaneous Repulsive Weighted HK model. After 100 simulations with $\varepsilon=0.1$ and 100 normal agents with a manipulative group with $K\in\{5j\}_{j=0}^{30}$ agents changing their opinions from $f(0)=-0.8$ to $f(t_{\Delta})=0.8$ in $t_{\Delta}\in\{5j\}_{j=1}^{30}$ iterations. The graphs show the averages of the mean opinions (left) and the average of final opinions (right).
Figure 23. Simultaneous Repulsive Weighted HK model. Simulation with the same parameters as in figure 22 except the confidence threshold, which is constant $\varepsilon=0.6$.
Observe that, for $\varepsilon=0.1$ and $t=t_{\Delta}$ iterations, the effect of the manipulative group’s action increases approximately to $\overline{\mathbf{x}}=0.35$ the average of the normal opinions when the values of iterations $t_{\Delta}$ and the number of manipulators $K$ are large (yellows in figure 22, top left). Recall that, for this simulation, we have chosen values of $K$ up to 1.5 times the number of normal population $N=100$. Bearing in mind that, without the influence of the manipulators, the opinions are grouped into two clusters of (approximately) the same size at $x=\pm 1$. Let us call $\mu$ the number of normal agents that pass from the group that converges to $x=-1$ to the group that converges to $x=1$. When a manipulative group acts obtaining the same two final clusters at $x=\pm 1$, we can compute the average
$\overline{\mathbf{x}}=\frac{(50+\mu)-(50-\mu)}{100}=\frac{\mu}{50}$
then if the average satisfies $0\leq\overline{\mathbf{x}}\leq 0.35$, then $0\leq\mu\leq 17$. This implies that a manipulative group with 150 agents is not able to modify the opinion of more than 17 normal agents (17 % of the total normal population).
The dispersion of opinions at $t=t_{\Delta}$ decreases as $t_{\Delta}$ increases, because when this value is small, the number of iterations is not large enough for the dynamics to stabilize (figure 22, top right). This stabilization can be reached in the range $t\in[t_{\Delta},1000]$ when the manipulative group behaves as stubborn.
If we compare with the results at $t=1000$, the final distributions consist of two big groups at $x=\pm 1$ and some isolated agents with moderated opinions. In addition, as in the attractive model, for small values of $t_{\Delta}$, the value of the mean moves in the opposite direction to the movement of the manipulator group (in dark blue in figure 22, bottom left) with a major effect reaching $x=-0.35$.
In the case of $\varepsilon=0.6$ (figure 23), we observe that the average of normal opinions at $t=t_{\Delta}$ is, approximately, in the range $0.20\leq\overline{\mathbf{x}}<$ 0.2 and the distribution of the normal opinion consists, mainly, in two clusters at $x=\pm 1$ for values $t_{\Delta}\geq 50$ (dark blue in figure 23, top right). When we consider $t=1000$ iterations, the final distribution is bimodal $x=\pm 1$ for any values of $t_{\Delta}$ and $K$.
3.5.3. Attractive–Repulsive Weighted HK model
We have seen in section 2.3 that the dynamics of the simultaneous Attractive–Repulsive Weighted HK model, for small values of $\varepsilon$, consists of a set of opinions oscillating in the central range of the opinion space, while extreme opinions $x=\pm 1$ can also group close normal agents together. Some central oscillating opinions can escape to converge to one of those extremal clusters.
With the intervention of manipulators, normal opinions outside the confidence interval of the manipulative group are moved away from this group. Because this interval is small, the opinions that the manipulative group can attract are few, and this attraction can be maintained throughout the iterations when the variation of the manipulative opinions is slow, that is, when $t_{\Delta}$ is large.
Thus, in the simulation with 100 normal agents, a manipulative group of 80 agents whose opinions vary from $f(0)=-0.8$ to $f(t_{\Delta})=0.8$ and a confidence threshold of $\varepsilon=0.1$, it is observed (cf. figure 24) that, as the iterations run to $t=t_{\Delta}$, the manipulative group presses the normal opinions with higher values towards $x=1$, making their convergence to this cluster faster. On the other hand, the range of opinions lower than that of the manipulators becomes wider and allows normal opinions to oscillate. This oscillation will be displaced downwards by the action of the repulsive mechanism when the manipulative group becomes stubborn. As $t_{\Delta}$ gets bigger, these oscillating opinions eventually converge to cluster $x=-1$ (cf. figure 24, center and right).
Figure 24. Simultaneous Attractive–Repulsive Weighted HK model. Simulation with 100 normal agents interacting with a group of 80 manipulators with opinions initial $f(0)=-0.8$ and final $f(t_{\Delta})=0.8$ reached in $t_{\Delta}\in\{20,40,60\}$ iterations. The confidence threshold is fixed at $\varepsilon=0.1$ constant.
When the confidence threshold $\varepsilon$ increases, the attractive mechanism also increases its influence while the repulsive decreases its own, allowing the manipulating group to attract a greater number of normal agents. This can be seen in figure 25, which corresponds to the simulation with the same parameters as in figure 24 except that we have fixed $\varepsilon=0.6$. In the three graphics of figure 25, we can observe that normal opinions are attracted by the manipulative group (dashed line) until the iteration $t=t_{\Delta}$ and, subsequently, when it becomes a stubborn group.
Figure 25. Simultaneous Attractive–Repulsive Weighted HK model: Simulation with the same parameters as in figure 24 except the confidence threshold, which is constant $\varepsilon=0.6$.
To study the dependence of the mean opinions on the number of manipulators $K$ and the number of iterations $t_{\Delta}$ during which the manipulative group varies its opinion, we will consider simulations for $\varepsilon\in\{0.1,0.6\}$, analogous to those in the previous section. Thus, we will set $N=100$ the number of normal agents, while the variables $t_{\Delta}$ and $K$ take values in the ranges $t_{\Delta}\in\{5j\}_{j=1}^{40}$ and $K\in\{5j\}_{j=0}^{40}$. In this case, we have chosen to extend the ranges of $t_{\Delta}$ and $K$ in order to better observe the changes in mean opinions.
For the confidence threshold $\varepsilon=0.1$, we can notice that the average of the mean opinions at iteration $t=t_{\Delta}$ (see figure 26 top left) reaches approximately the value $x\simeq 0.5$ if $K\simeq 200$ and $t_{\Delta}\simeq 200$, but remains close to $x=0$ when $K\lesssim 80$ (in blue colors). The mean opinion will increase as the value of $K$ increases, and specially from $K\simeq 80$, and favouring this increase in the mean opinion, the large values of $t_{\Delta}$ over the small ones.
It is important to note that this fact implies that, in order for the manipulative group to have a significant influence on the average opinion, the size of the group must be greater than 80% of the size of the normal population, unlike what happens in the HK and DW models, in which a smaller manipulative group can have a much more important influence.
Again, because of the tendency to polarization in a bimodal distribution at $x=\pm 1$, even with the existence of moderate oscillating opinions, the average of the standard deviations is large and the number of clusters approaches 2 (in dark blue in figure 26 at the top right) when the values of $t_{\Delta}$ cease to be small.
Figure 26. Simultaneous Attractive–Repulsive Weighted HK model. After 100 simulations with $\varepsilon=0.1$ and 100 normal agents with a manipulative group with $K\in\{5j\}_{j=0}^{40}$ agents changing their opinions from $f(0)=-0.8$ to $f(t_{\Delta})=0.8$ in $t_{\Delta}\in\{5j\}_{j=1}^{40}$ iterations. The graphs show the averages of the mean opinions (left), the standard deviation (centre) and the number of final opinions (right).
In the bottom row of figure 26, we show the results after $t=500$ iterations in such a way that the opinion of the manipulative group is $f(t)=0.8$ for all $t\in\{t_{\Delta},\ldots,500\}$, that is, the manipulators become stubborn agents. In this case, oscillating opinions (if any) tend to converge to one of the extreme clusters (or may remain oscillating), stabilizing the average values studied.
Thus, when $t_{\Delta}$ is small, the averages of the mean opinions move in the opposite direction to the opinion of the manipulative group, even reaching $x\simeq-0.25$. Again, this is because the manipulative group changes its opinion quickly and before the normal opinions converge to one of the clusters, and since the confidence interval is small, the repulsive mechanism pushes the normal opinions in the opposite direction to that of the manipulators (see dark blue region in the figure 26 bottom left). The stabilization of the values is also observed by the reduction in the final number of clusters, which are between 2 and 9 (cf. figure 26 bottom right).
When we set the confidence threshold at $\varepsilon=0.6$, the attractive mechanism has greater influence and the manipulative group is able to attract a greater number of normal opinions. After $t=t_{\Delta}$ iterations (see top row of figure 27), we can observe that the highest average of the mean opinions is reached for values $(t_{\Delta},K)\simeq(40,200)$ (in yellow in figure 27 top left). For greater values of $t_{\Delta}$, the manipulative group needs more iterations to reach its final opinion. Thus, normal opinions in the confidence interval of the manipulative group are more likely to be repelled by other agents while they follow the manipulative group (seen in figure 25).
It is also observed, even more clearly than in the case of $\varepsilon=0.1$, that when $K\lesssim 60$, the influence of the manipulating group is small (note clearly delimited the dark blue region in figure 27 above left). The behaviour of the average of the final number of clusters is similar to the $\varepsilon=0.1$ case, in which, only for small values of $t_{\Delta}$, there are a large number of opinions that are not yet clustered.
When we consider the dynamics after $t=500$ iterations, we observe that the average of the mean opinions remains similar except for small values of $t_{\Delta}$, for which the effect of the shift in the opposite direction to that of the manipulators’ opinion is repeated. As for the number of clusters, the graph at the bottom right of figure 27 shows, in dark blue, a region in which the final distribution converges to the extreme values $x=\pm 1$, except in the yellow region, which corresponds to a distribution with a large number of opinions that oscillate attracted by the cluster $x=1$ and the manipulative group (now stubborn) at $x=0.8$ as seen in figure 25 left. This region of oscillating opinions is clearly delimited as figure shows.
Figure 27. Simultaneous Attractive–Repulsive Weighted HK model. Simulation with the same parameters as in figure 26 except the confidence threshold, which is constant $\varepsilon=0.6$.
4. Conclusions and discussion
We have studied the opinion dynamics in various models of bounded confidence under the influence of one group of manipulative agents. This type of agent, as a generalization or extension of stubborn agents, allows them to have influence within the union of the confidence intervals corresponding to their range of opinions. Thus, strategies can be established to modify the natural dynamics of the model (due only to normal agents). This modification can cause opinions acceptable to society of normal agents to become extreme or vice versa. The range of opinions can be moved as well as expand or contract in the space of opinions, thus reminiscent of the Overton window theory. Therefore, manipulators may attempt to modify the Overton window to form a new framework of acceptable opinions.
In this article, we have described the strategies of one single manipulative group that varies its opinions linearly over a wide range of opinions, and then becomes a stubborn group to the point of convergence or stabilization of the dynamics. The number of manipulators in the group, as well as the rate of variation in their opinions, have an unequal influence on the different models studied, with the different confidence thresholds. Due to the arbitrariness of choosing a canonical definition for the Overton window, we consider that the mean as the central point and the standard deviation as the amplitude can be used to determine it. We have shown examples of simulations of opinion dynamics that allow us to interpret the results of the change in the Overton window, using mean values and standard deviations, based on the variables that define the manipulative group. These changes in values resulting from dynamics are quantified with a group of manipulators that move in a wide and fixed range for the different models. In the HK and DW models, because of their tendency to central clustering, a manipulative group with a sufficient number of agents and/or whose change of opinion is sufficiently slow, can attract a large number of normal agents and take them with them to any point in the opinion space. Moreover, for the simultaneous HK model, the conditions of these variables have been fully determined so that manipulators attract the opinions of another group of normal agents. In Weighted HK models, a larger number of manipulators is needed to influence the opinion of normal agents, and rapid changes in the opinion of manipulators could have the opposite effect on normal agents.
There are still pending issues that will be studied in future articles, among them, the competition between various groups of manipulators according to their size and the trajectories of their opinions, the finding of new models that adapt to the movement of the Overton window placing it in the central point of the space of opinions. In this sense, the density function $\phi_{i}(t)=1-|x_{i}(t)|$ in the iteration of the Weighted HK model can be modified so that its mean (or median) in each iteration matches the mean (or median) of the distribution of normal opinions.
Acknowledgements
The author would like to thank Gemma Rioja for her invaluable suggestions and comments that have substantially improved the text of this article.
Data availability
No new data were created or analysed during this study. Data sharing is not applicable to this article.
All the simulations and graphs in this article have been made using the R2020b version of Matlab.